# On local Galois deformation rings: generalised reductive groups

### v1

@@ 7. Lafforgue’s $G$-pseudocharacters
In his work on the local Langlands correspondence for function fields [36] Vincent Lafforgue introduced a notion of $G$-pseudocharacter for a reductive group $G$. If $G=\operatorname{GL}_{d}$ then Emerson and Morel [25] have shown that $\operatorname{GL}_{d}$-pseudocharacter after Lafforgue is equivalent to a determinant law after Chenevier [16]. If $d!$ is invertible in the base ring then the trace of representation determines its $\operatorname{GL}_{d}$-pseudocharacter. In the arithmetic setting this was first used by Wiles [58] and Taylor [55].
@@ 7.1. Definition of $G$-pseudocharacter
The definition of $G$-pseudocharacter we use is a slight modification of Lafforgue’s original definition [36, Section 11], in that we work over the base ring $\mathcal{O}$ and allow arbitrary disconnected generalised reductive group schemes for $G$.
@@
When $\mathcal{O}^{\prime}$ is a commutative $\mathcal{O}$-algebra the natural maps $\mathcal{O}[G^{n}]^{G^{0}}\to\mathcal{O}^{\prime}[G^{n}]^{G^{0}}$ induce a map $\mathrm{PC}_{G_{\mathcal{O}^{\prime}}}^{\Gamma}(A)\to\mathrm{PC}_{G}^{\Gamma}(A)$. This map is always a bijection. When $\mathcal{O}^{\prime}$ is flat over $\mathcal{O}$ the map $\mathcal{O}[G^{n}]^{G^{0}}\otimes_{\mathcal{O}}\mathcal{O}^{\prime}\to\mathcal{O}^{\prime}[G^{n}]^{G^{0}}$ is an isomorphism by the universal coefficient theorem in rational cohomology. The case of a general $\mathcal{O}$-algebra is treated in [44, Proposition 3.22]. We will use this fact whenever we need to extend the base ring $\mathcal{O}$.
@@ 7.2. Deformations of $G$-pseudocharacters
We adopt the deformation theoretic setup of subsection 5.2. Let $\kappa$ be either a finite or a local field, which is an $\mathcal{O}$-algebra, equipped with its natural topology. Let $\Lambda$ be the coefficient ring for $\kappa$ and $\mathfrak{A}_{\Lambda}$ is the category of local artinian $\Lambda$-algebras with residue field $\kappa$, equipped with the natural topology.
@@ 7.3. Comparison of $X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ with $X^{\mathrm{ps}}_{G}$
In this section we assume that $\Gamma$ is a profinite topologically finitely generated group. It has been shown by the second author in [44, Theorem A] that $R^{\mathrm{ps}}_{G}$ is a complete noetherian local $\Lambda$-algebra.
@@
that attaches to a representation $\rho$ its $G$-pseudocharacter $\Theta_{\rho}$. By [44, Theorem 3.20] $\mathrm{PC}_{G}^{\Gamma}$ is representable by an affine scheme, so $\vartheta:X^{\mathrm{gen},\tau}_{G}\to\mathrm{PC}_{G}^{\Gamma}$ factors over $\theta:X^{\mathrm{git},\tau}_{G}\to\mathrm{PC}_{G}^{\Gamma}$.
@@
By [54, Tag 04DF], it suffices to show, that $\nu$ is finite, universally injective and surjective. By 6.4 $R_{G,\overline{\rho}^{\mathrm{ss}}}^{\mathrm{git},\tau}$ is a finite $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-module. From the embedding $\tau:G\to\operatorname{GL}_{d}$ we obtain a factorisation $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\to R^{\mathrm{ps}}_{G}\to R^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$. In particular, $R^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ is a finite $R^{\mathrm{ps}}_{G}$-module.
@@
The closed point of $X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ is mapped to the closed point of $X^{\mathrm{ps}}_{G}$, so we only need to show surjectivity of $\dot{\nu}:U^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}:=X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}\setminus\{*\}\to U^{\mathrm{ps}}_{G}:=X^{\mathrm{ps}}_{G}\setminus\{*\}$. By [54, Tag 02J6] $U^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ and $U^{\mathrm{ps}}_{G}$ are Jacobson. By Chevalley’s theorem [54, Tag 054K] the image of $\dot{\nu}$ is locally constructible and thus also its complement. Therefore for surjectivity of $\nu$ it is sufficient to show, that the closed points of $U^{\mathrm{ps}}_{G}$ are all contained in the image of $\dot{\nu}$.
@@
Let $y\in U^{\mathrm{ps}}_{G}$ be a closed point. By [8, Lemma 3.17] $\kappa(y)$ is a local field. By [44, Theorem 5.7], $R^{\mathrm{ps}}_{G}$ is noetherian, and thus $\kappa(y)$ with its natural topology is a topological $R^{\mathrm{ps}}_{G}$-algebra and the image of $y:R^{\mathrm{ps}}_{G}\rightarrow\kappa(y)$ is contained in the ring of integers $\mathcal{O}_{\kappa(y)}$ by [8, Lemma 3.17]. We obtain a continuous $G$-pseudocharacter $\Theta_{y}\in\mathrm{cPC}_{G}^{\Gamma}(\mathcal{O}_{\kappa(y)})$, which is a deformation of $\Theta_{\overline{\rho}}$ to $\mathcal{O}_{\kappa(y)}$. By the continuous reconstruction theorem [44, Theorem 3.8], there is a continuous representation $\rho^{\prime}:\Gamma\to G(\overline{\kappa})$ with $\Theta_{\rho^{\prime}}=\Theta_{y}$, where $\overline{\kappa}$ is an algebraic closure of $\kappa(y)$. Lemma 5.12 implies that $\rho^{\prime}\in X^{\mathrm{gen},\tau}_{G}(\overline{\kappa})$. The $G(\overline{\kappa})$-conjugacy class of $\rho^{\prime}$ defines a point $y^{\prime}\in X^{\mathrm{git},\tau}_{G}(\overline{\kappa})$ which maps to $y$. We have to show that $y^{\prime}$ lies on the same connected component of $X^{\mathrm{git},\tau}_{G}$ as $\overline{\rho}^{\mathrm{ss}}$. Since $X^{\mathrm{gen},\tau}_{G}$ is of finite type over $X^{\mathrm{ps}}_{\operatorname{GL}_{d}}$, $\rho^{\prime}$ takes values in $G(\kappa^{\prime})$ for some finite extension of $\kappa(y)$ in $\overline{\kappa}$. Proposition 6.6 implies that after conjugating $\rho^{\prime}$ with an element of $G^{0}(\overline{\kappa})$ we may assume that $\rho^{\prime}$ takes values in $G(\mathcal{O}_{\overline{\kappa}})$. Let $\overline{\rho}^{\prime}$ be the composition of $\rho^{\prime}$ with the reduction map $G(\mathcal{O}_{\overline{\kappa}})\rightarrow G(\overline{k})$. Then $\Theta_{\overline{\rho}^{\prime}}=\Theta_{y}\otimes_{\mathcal{O}_{\kappa(y)}}\overline{k}=\Theta_{\overline{\rho}^{\mathrm{ss}}}$. Uniqueness part of the reconstruction theorem implies that $G$-semisimplification of $\overline{\rho}^{\prime}$ is equal to $\overline{\rho}^{\mathrm{ss}}$, and Lemma 6.12 implies that $y^{\prime}\in X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$.
@@
For universal injectivity it suffices by [54, Tag 01S4] to show that the diagonal morphism $X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}\to X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}\times_{X^{\mathrm{ps}}_{G}}X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ is surjective. Let $x\in X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}\times_{X^{\mathrm{ps}}_{G}}X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$. We obtain by 6.7 and 6.8 two $G$-semisimple representations $\rho_{x},\rho_{x}^{\prime}\in X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}(\overline{\kappa(x)})$, such that their associated $G$-pseudocharacters coincide. This means by the uniqueness part of the reconstruction theorem [44, Theorem 3.7] and 6.10, their image in $X^{\mathrm{git}}_{G}(\overline{\kappa(x)})$ coincides and therefore $x$ is in the image of the diagonal. ∎
@@
The bottom map is surjective, since $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}[1/p][\operatorname{Rep}_{G}^{\mathcal{F}_{N},\square}]\to A^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]$ is surjective as explained above and taking $G^{0}$-invariants is exact in characteristic 00. Since the diagram commutes the right vertical arrow is surjective, and we obtain the claim. ∎
@@ 8. Functoriality
We assume in this section that $\Gamma$ is topologically finitely generated. This assumption ensures that $R^{\mathrm{ps}}_{G}$ is noetherian. If $\Gamma$ is only assumed to satisfy Mazur’s $p$-finiteness condition then results of this section can be reformulated by replacing $\Gamma$ by a topologically finitely generated quotient as in Lemma 5.5. We note that $\Gamma_{F}$ is topologically finitely generated by [34, Satz 3.6].
@@ 8.1. Independence of $\tau$ and functoriality in $G$
Conversely, let $\rho\in X^{\mathrm{gen},\tau}_{G}(A)$ satisfy $\Theta_{\rho}=\Theta^{u}\otimes_{R^{\mathrm{ps}}_{G}}A$. By factoring $\rho$ as $\Gamma\to G(A^{\mathrm{gen},\tau}_{G})\to G(A)$ over the universal representation on $A^{\mathrm{gen},\tau}_{G}$ and using that $A^{\mathrm{gen},\tau}_{G}$ is of finite type over $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$, we see that $\rho$ actually takes values in a finitely generated $R^{\mathrm{ps}}_{G}$-subalgebra $A^{\prime}\subseteq A$. We want to show, that the image of $\operatorname{Spec}(A^{\prime})\to X^{\mathrm{gen},\tau}_{G}$ is entirely contained in the component $X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$, which amounts to showing that an idempotent $e\in A^{\mathrm{gen},\tau}_{G}$ associated to a component $X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}_{2}}$ associated with a residual representation $\overline{\rho}^{\mathrm{ss}}_{2}$ which is non-conjugate to $\overline{\rho}^{\mathrm{ss}}$ acts as zero on $A^{\prime}$. By contradiction, we assume that $eA^{\prime}$ is nonzero. As $A^{\prime}$ is of finite type over $R^{\mathrm{ps}}_{G}$, there is a dimension $\leq 1$ prime ideal $\mathfrak{p}\in\operatorname{Spec}(eA^{\prime})$ with residue field $\kappa(\mathfrak{p})$ finite or local by [8, Lemma 3.17]. By 5.12, we know that the representation $\rho_{\mathfrak{p}}:\Gamma\to G(\overline{\kappa(\mathfrak{p})})$ is continuous and that $\Theta_{\rho_{\mathfrak{p}}}=\Theta^{u}\otimes_{R^{\mathrm{ps}}_{G}}\overline{\kappa(\mathfrak{p})}$. If $\kappa(\mathfrak{p})$ is a finite field then we deduce that the $G$-semisimplification of $\rho_{\mathfrak{p}}$ is $\overline{\rho}^{\mathrm{ss}}$, which yields a contradiction to the choice of $e$ via Lemma 6.13. If $\kappa(\mathfrak{p})$ is a local field, by 6.6 the image of $\rho_{\mathfrak{p}}$ lies up to conjugation in the ring of integers of a finite extension of $\kappa(\mathfrak{p})$. The $G$-semisimplification of the reduction modulo the maximal ideal is $\overline{\rho}^{\mathrm{ss}}$, since $\Theta_{\rho_{\mathfrak{p}}}$ specialises $\Theta^{u}$, which yields a contradiction to the choice of $e$ via Lemma 6.13. We conclude, that $\rho\in X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}(A^{\prime})\subseteq X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}(A)$. ∎
@@
If $\rho\in X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}(A)$ then $\Theta_{\rho}=\Theta^{u}_{|A}$ by Lemma 8.1. Moreover, $\tau\circ\rho$ is $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-condensed by Lemma 5.10 (1). Lemmas 4.16 and 4.5 implies that $\rho$ is $R^{\mathrm{ps}}_{G}$-condensed.
If $\rho:\Gamma\rightarrow G(A)$ satisfies the conditions of part (2) then $\det\circ(\tau\circ\rho)^{\mathrm{lin}}=D^{u}_{|A}$ and $\tau\circ\rho$ is $R^{\mathrm{ps}}_{G}$-condensed. Thus $\rho$ satisfies part (4) of Lemma 5.2 with $R=R^{\mathrm{ps}}_{G}$ and thus $\rho\in X^{\mathrm{gen},\tau}_{G}(A)$ by Lemma 5.10 (1). Since $\Theta_{\rho}=\Theta^{u}_{|A}$ Lemma 8.1 implies that $\rho\in X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}(A)$. ∎
@@
Let $\varphi:G\to H$ be a homomorphism of generalised reductive $\mathcal{O}$-group schemes. Then there is a natural map of $R^{\mathrm{ps}}_{H}$-schemes $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}\to X^{\mathrm{gen}}_{H,\varphi\circ\overline{\rho}^{\mathrm{ss}}}$, such that for every $R^{\mathrm{ps}}_{G}$-algebra $A$, the map $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}(A)\to X^{\mathrm{gen}}_{H,\varphi\circ\overline{\rho}^{\mathrm{ss}}}(A)$ is given by $\rho\mapsto\varphi\circ\rho$.

### v2

@@ 7. Lafforgue’s $G$-pseudocharacters
In his work on the local Langlands correspondence for function fields [37] Vincent Lafforgue introduced a notion of $G$-pseudocharacter for a reductive group $G$. If $G=\operatorname{GL}_{d}$ then Emerson and Morel [25] have shown that a $\operatorname{GL}_{d}$-pseudocharacter after Lafforgue is equivalent to a determinant law after Chenevier [15]. If $d!$ is invertible in the base ring then the trace of representation valued in $\operatorname{GL}_{d}$ determines its $\operatorname{GL}_{d}$-pseudocharacter. In the arithmetic setting this was first used by Wiles [63] and Taylor [60].
@@ 7.1. Definition of $G$-pseudocharacter
The definition of $G$-pseudocharacter we use is a slight modification of Lafforgue’s original definition [37, Section 11], in that we work over the base ring $\mathcal{O}$ and allow arbitrary (disconnected) generalised reductive group schemes for $G$.
@@
When $\mathcal{O}^{\prime}$ is a commutative $\mathcal{O}$-algebra and $A$ is a commutative $\mathcal{O}^{\prime}$-algebra the natural maps $\mathcal{O}[G^{n}]^{G^{0}}\to\mathcal{O}^{\prime}[G^{n}]^{G^{0}}$ induce a map $\mathrm{PC}_{G_{\mathcal{O}^{\prime}}}^{\Gamma}(A)\to\mathrm{PC}_{G}^{\Gamma}(A)$. This map is always a bijection, as we will now explain. When $\mathcal{O}^{\prime}$ is flat over $\mathcal{O}$ the map $\mathcal{O}[G^{n}]^{G^{0}}\otimes_{\mathcal{O}}\mathcal{O}^{\prime}\to\mathcal{O}^{\prime}[G^{n}]^{G^{0}}$ is an isomorphism by the universal coefficient theorem in rational cohomology, see [35, Proposition I.4.18 (a)]. The case of a general $\mathcal{O}$-algebra is treated in [50, Proposition 3.22]. We will use this fact whenever we need to extend the base ring $\mathcal{O}$.
@@ 7.2. Deformations of $G$-pseudocharacters
We adopt the deformation theoretic setup of subsection 5.2. Let $\kappa$ be either a finite or a local field, which is an $\mathcal{O}$-algebra, equipped with its natural topology. Let $\Lambda$ be the coefficient ring for $\kappa$ and let $\mathfrak{A}_{\Lambda}$ be the category of local artinian $\Lambda$-algebras with residue field $\kappa$, equipped with the natural topology.
@@ 7.3. Comparison of $X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ with $X^{\mathrm{ps}}_{G}$
In this section we assume that $\Gamma$ is a profinite topologically finitely generated group. It has been shown by the second author in [50, Theorem A] that $R^{\mathrm{ps}}_{G}$ is a complete noetherian local $\Lambda$-algebra.
@@
that attaches to a representation $\rho$ its $G$-pseudocharacter $\Theta_{\rho}$. By [50, Theorem 3.20] $\mathrm{PC}_{G}^{\Gamma}$ is representable by an affine scheme, so $\vartheta:X^{\mathrm{gen},\tau}_{G}\to\mathrm{PC}_{G}^{\Gamma}$ factors over $\theta:X^{\mathrm{git},\tau}_{G}\to\mathrm{PC}_{G}^{\Gamma}$.
@@
By [59, Tag 04DF], it suffices to show, that $\nu$ is finite, universally injective and surjective. By 6.4 $R_{G,\overline{\rho}^{\mathrm{ss}}}^{\mathrm{git},\tau}$ is a finite $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-module. From the embedding $\tau:G\to\operatorname{GL}_{d}$ we obtain a factorisation $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\to R^{\mathrm{ps}}_{G}\to R^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$. In particular, $R^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ is a finite $R^{\mathrm{ps}}_{G}$-module.
@@
The closed point of $X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ is mapped to the closed point of $X^{\mathrm{ps}}_{G}$, so we only need to show surjectivity of $\dot{\nu}:U^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}:=X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}\setminus\{*\}\to U^{\mathrm{ps}}_{G}:=X^{\mathrm{ps}}_{G}\setminus\{*\}$. By [59, Tag 02J6] $U^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ and $U^{\mathrm{ps}}_{G}$ are Jacobson. By Chevalley’s theorem [59, Tag 054K] the image of $\dot{\nu}$ is locally constructible and thus also its complement. Therefore for surjectivity of $\nu$ it is sufficient to show, that the closed points of $U^{\mathrm{ps}}_{G}$ are all contained in the image of $\dot{\nu}$.
@@
Let $y\in U^{\mathrm{ps}}_{G}$ be a closed point. By [6, Lemma 3.17] $\kappa(y)$ is a local field. By [50, Theorem 5.7], $R^{\mathrm{ps}}_{G}$ is noetherian, and thus $\kappa(y)$ with its natural topology is a topological $R^{\mathrm{ps}}_{G}$-algebra and the image of $y:R^{\mathrm{ps}}_{G}\rightarrow\kappa(y)$ is contained in the ring of integers $\mathcal{O}_{\kappa(y)}$ by [6, Lemma 3.17]. We obtain a continuous $G$-pseudocharacter $\Theta_{y}\in\mathrm{cPC}_{G}^{\Gamma}(\mathcal{O}_{\kappa(y)})$, which is a deformation of $\Theta_{\overline{\rho}}$ to $\mathcal{O}_{\kappa(y)}$. By the continuous reconstruction theorem [50, Theorem 3.8], there is a continuous representation $\rho^{\prime}:\Gamma\to G(\overline{\kappa})$ with $\Theta_{\rho^{\prime}}=\Theta_{y}$, where $\overline{\kappa}$ is an algebraic closure of $\kappa(y)$. Lemma 5.12 implies that $\rho^{\prime}\in X^{\mathrm{gen},\tau}_{G}(\overline{\kappa})$. The $G(\overline{\kappa})$-conjugacy class of $\rho^{\prime}$ defines a point $y^{\prime}\in X^{\mathrm{git},\tau}_{G}(\overline{\kappa})$ which maps to $y$. We have to show that $y^{\prime}$ lies on the same connected component of $X^{\mathrm{git},\tau}_{G}$ as $\overline{\rho}^{\mathrm{ss}}$. Since $X^{\mathrm{gen},\tau}_{G}$ is of finite type over $X^{\mathrm{ps}}_{\operatorname{GL}_{d}}$, $\rho^{\prime}$ takes values in $G(\kappa^{\prime})$ for some finite extension of $\kappa(y)$ in $\overline{\kappa}$. Proposition 6.6 implies that after conjugating $\rho^{\prime}$ with an element of $G^{0}(\overline{\kappa})$ we may assume that $\rho^{\prime}$ takes values in $G(\mathcal{O}_{\overline{\kappa}})$. Let $\overline{\rho}^{\prime}$ be the composition of $\rho^{\prime}$ with the reduction map $G(\mathcal{O}_{\overline{\kappa}})\rightarrow G(\overline{k})$. Then $\Theta_{\overline{\rho}^{\prime}}=\Theta_{y}\otimes_{\mathcal{O}_{\kappa(y)}}\overline{k}=\Theta_{\overline{\rho}^{\mathrm{ss}}}$. The uniqueness part of the reconstruction theorem [50, Theorem 3.7] implies that $G$-semisimplification of $\overline{\rho}^{\prime}$ is equal to $\overline{\rho}^{\mathrm{ss}}$, and Lemma 6.12 implies that $y^{\prime}\in X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$.
@@
For universal injectivity it suffices by [59, Tag 01S4] to show that the diagonal morphism $X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}\to X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}\times_{X^{\mathrm{ps}}_{G}}X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ is surjective. Let $x\in X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}\times_{X^{\mathrm{ps}}_{G}}X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$. We obtain by 6.7 and 6.8 two $G$-semisimple representations $\rho_{x},\rho_{x}^{\prime}\in X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}(\overline{\kappa(x)})$, such that their associated $G$-pseudocharacters coincide. This means by the uniqueness part of the reconstruction theorem and 6.10, their image in $X^{\mathrm{git}}_{G}(\overline{\kappa(x)})$ coincides and therefore $x$ is in the image of the diagonal. ∎
@@
The bottom map is surjective, since $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}[1/p][\operatorname{Rep}_{G}^{\mathcal{F}_{N},\square}]\to A^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]$ is surjective as explained above and taking $G^{0}$-invariants is exact in characteristic 0. Since the diagram commutes the right vertical arrow is surjective, and we obtain the claim. ∎
@@
Remark 7.7.
In [47], building on the results of this paper, we show that $\nu\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}$ is an isomorphism.
@@ 8. Functoriality
We assume in this section that $\Gamma$ is topologically finitely generated. This assumption ensures that $R^{\mathrm{ps}}_{G}$ is noetherian. If $\Gamma$ is only assumed to satisfy Mazur’s $p$-finiteness condition [^7] In [46], using the results of this paper, we show that if $\Gamma_{F}$ satisfies Mazur’s $p$-finiteness condition then $R^{\mathrm{ps}}_{G}$ is noetherian and the results of this section still hold in this more general setting. then results of this section can be reformulated by replacing $\Gamma$ by a topologically finitely generated quotient as in Lemma 5.5. We note that $\Gamma_{F}$ is topologically finitely generated by [34, Satz 3.6].
@@ 8.1. Independence of $\tau$ and functoriality in $G$
Conversely, let $\rho\in X^{\mathrm{gen},\tau}_{G}(A)$ satisfy $\Theta_{\rho}=\Theta^{u}\otimes_{R^{\mathrm{ps}}_{G}}A$. By factoring $\rho$ as $\Gamma\to G(A^{\mathrm{gen},\tau}_{G})\to G(A)$ over the universal representation on $A^{\mathrm{gen},\tau}_{G}$ and using that $A^{\mathrm{gen},\tau}_{G}$ is of finite type over $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$, we see that $\rho$ actually takes values in a finitely generated $R^{\mathrm{ps}}_{G}$-subalgebra $A^{\prime}\subseteq A$. We want to show, that the image of $\operatorname{Spec}(A^{\prime})\to X^{\mathrm{gen},\tau}_{G}$ is entirely contained in the component $X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$, which amounts to showing that an idempotent $e\in A^{\mathrm{gen},\tau}_{G}$ associated to a component $X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}_{2}}$ associated with a residual representation $\overline{\rho}^{\mathrm{ss}}_{2}$ which is non-conjugate to $\overline{\rho}^{\mathrm{ss}}$ acts as zero on $A^{\prime}$. By contradiction, we assume that $eA^{\prime}$ is nonzero. As $A^{\prime}$ is of finite type over $R^{\mathrm{ps}}_{G}$, there is a dimension $\leq 1$ prime ideal $\mathfrak{p}\in\operatorname{Spec}(eA^{\prime})$ with residue field $\kappa(\mathfrak{p})$ finite or local by [6, Lemma 3.17]. By 5.12, we know that the representation $\rho_{\mathfrak{p}}:\Gamma\to G(\overline{\kappa(\mathfrak{p})})$ is continuous and that $\Theta_{\rho_{\mathfrak{p}}}=\Theta^{u}\otimes_{R^{\mathrm{ps}}_{G}}\overline{\kappa(\mathfrak{p})}$. If $\kappa(\mathfrak{p})$ is a finite field then we deduce that the $G$-semisimplification of $\rho_{\mathfrak{p}}$ is $\overline{\rho}^{\mathrm{ss}}$, which yields a contradiction to the choice of $e$ via Lemma 6.13. If $\kappa(\mathfrak{p})$ is a local field, by 6.6 the image of $\rho_{\mathfrak{p}}$ lies up to conjugation in the ring of integers of a finite extension of $\kappa(\mathfrak{p})$. The $G$-semisimplification of the reduction modulo the maximal ideal is $\overline{\rho}^{\mathrm{ss}}$, since $\Theta_{\rho_{\mathfrak{p}}}$ specialises $\Theta^{u}$, which yields a contradiction to the choice of $e$ via Lemma 6.13. We conclude, that $\rho\in X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}(A^{\prime})\subseteq X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}(A)$. ∎
@@
If $\rho\in X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}(A)$ then $\Theta_{\rho}=\Theta^{u}_{|A}$ by Lemma 8.1. Moreover, $\tau\circ\rho$ is $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-condensed by Lemma 5.10 (1). Lemmas 4.16 and 4.5 imply that $\rho$ is $R^{\mathrm{ps}}_{G}$-condensed.
If $\rho:\Gamma\rightarrow G(A)$ satisfies the conditions of part (2) then $\det\circ(\tau\circ\rho)^{\mathrm{lin}}=D^{u}_{|A}$ and $\tau\circ\rho$ is $R^{\mathrm{ps}}_{G}$-condensed. Thus $\rho$ satisfies part (4) of Lemma 5.2 with $R=R^{\mathrm{ps}}_{G}$ and thus $\rho\in X^{\mathrm{gen},\tau}_{G}(A)$ by Lemma 5.10 (1). Since $\Theta_{\rho}=\Theta^{u}_{|A}$, Lemma 8.1 implies that $\rho\in X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}(A)$. ∎
@@
Let $\varphi:G\to H$ be a homomorphism of generalised reductive $\mathcal{O}$-group schemes. Then there is a natural map [^8] As pointed out by referee the target of the map should be denoted by $X^{\mathrm{gen}}_{H,(\varphi\circ\overline{\rho}^{\mathrm{ss}})^{\mathrm{ss}}}$, as $\varphi\circ\overline{\rho}^{\mathrm{ss}}$ need not be $H$-semisimple in general. For ease of reading, we will continue to use this abuse of notation. of $R^{\mathrm{ps}}_{H}$-schemes $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}\to X^{\mathrm{gen}}_{H,\varphi\circ\overline{\rho}^{\mathrm{ss}}}$, such that for every $R^{\mathrm{ps}}_{G}$-algebra $A$, the map $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}(A)\to X^{\mathrm{gen}}_{H,\varphi\circ\overline{\rho}^{\mathrm{ss}}}(A)$ is given by $\rho\mapsto\varphi\circ\rho$.